OpenStudy (anonymous):

HElP!!! I need help with some multi step equations with variables on both sides. I will give medals and fan people! :-)

OpenStudy (jessiegonzales):

problem?

OpenStudy (anonymous):

.75x=0.25x+12

OpenStudy (anonymous):



OpenStudy (anonymous):

@Astrophysics @amberosales @amirreza1870 @americansniper12 @ayeshaafzal221 @anaaxcx @barreraA @barbdewysveck1 @brian16 @BlooooooBlahBLIIIIK @baewolfstar @Bubwros @chmvijay @CallMeKiki @CloverRacer @CherokeeGutierrez @CelestialTime @cyploveawesome @Donblue @DALLINATOR720 @destinymontes @DurableToaster @dehelloo @denise19 @EwikaEnding @Frostbite @fallengoddess @freemap @ganeshie8 @geny55 @GuitarNerd2015 @Gonin12 @horsegirl27 @HELPMEPLZ!! @HoneyLemon @hannah120899 @Hano @hxj9314 @inkyvoyd @is3535 @Irule4ever @Indonique @iambatman @jim_thompson5910 @jazzyspazzy @jakelee @jfernandes @jonny3909 @paki

OpenStudy (lynfran):

are we solving for "x"..?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Help?

OpenStudy (lynfran):

then 0.75x=0.25x+12 0= -0.75x+0.25x+12 0=-x/2+12 x/2=12 now multiply both sides of the equation by 2

OpenStudy (anonymous):

Thank you

OpenStudy (lynfran):

welcome